# Average-Case Local Computation Algorithms

### v1

@@
Abstract
We initiate the study of Local Computation Algorithms on average case inputs. In the Local Computation Algorithm (LCA) model, we are given probe access to a huge graph, and asked to answer membership queries about some combinatorial structure on the graph, answering each query with sublinear work.
For instance, an LCA for the $k$-spanner problem gives access to a sparse subgraph $H\subseteq G$ that preserves distances up to a factor of $k$. We build simple LCAs for this problem assuming the input graph is drawn from the well-studied Erdős-Rényi and Preferential Attachment graph models. In both cases, our spanners achieve size and stretch tradeoffs that are impossible to achieve for general graphs, while having dramatically lower query complexity than worst-case LCAs.
Our second result investigates the intersection of LCAs with Local Access Generators (LAGs). Local Access Generators provide efficient query access to a random object, for instance an Erdős-Rényi random graph. We explore the natural problem of generating a random graph together with a combinatorial structure on it. We show that this combination can be easier to solve than focusing on each problem by itself, by building a fast, simple algorithm that provides access to an Erdős-Rényi random graph together with a maximal independent set.
1 Introduction
When computing on a very large object, it can be important to find fast algorithms which answer user queries to the solution, while neither considering the whole input, nor computing the full output solution. In the local computation model [RTVX11, ARVX12], we are given probe access to a large object, such as a graph, and receive queries about some combinatorial structure on the graph. For example, in the case of the maximal independent set (MIS) problem, the queries ask whether node $u$ is in the maximal independent set. We desire Local Computation Algorithms (LCAs) that can quickly answer such queries while making very few probes to the graph. Moreover, we require the answers returned by the algorithm on different queries to be consistent with a fixed global structure. For the MIS example, the LCA must answer all possible queries according to a single maximal independent set. This consistency requirement is challenging since typically, we require the LCA to be memoryless – it does not store information about its previous answers. The only allowed coordination is in the form of a shared random string. Formally, the definition of LCA is as follows.
@@ 1 Introduction
A Local Computation Algorithm (LCA) for a problem $\Pi$ is an oracle $\mathcal{A}$ with the following properties. $\mathcal{A}$ is given probe access to the input $G$, a sequence of random bits $\vec{r}$ and sub-linear local memory. For any use query $q$ in a family of admissible queries to the output, $\mathcal{A}$ must use only its oracle access to $G$ (which we refer to as probes to $G$), random bits $\vec{r}$, and local memory to answer the query $q$. After answering the query, $\mathcal{A}$ erases its local memory (including the query $q$ and its response). All the responses to queries given by $\mathcal{A}$ must be consistent with a single valid solution $X$ to the specified computation problem on input $G$.
Thus, independent copies of $\mathcal{A}$ with the same $G$ and the same random bits $\vec{r}$, must agree on a single solution $X$ that is consistent with all answered queries, regardless of the order and content of queries that were actually asked. We generally desire LCAs to succeed with high probability over the random string $\vec{r}$, and we wish to minimize the time and probe complexity of answering each query.
There has been extensive work both obtaining sublinear time LCAs for a variety of natural problems, and further optimizing the per-query work to be as fast as possible. For example, on bounded degree graphs, there are LCAs with polylogarithmic query complexity for spanners [LRR16, LL18, LRR20, PRVY19, ACLP23, BF24], maximal independent set (MIS) [Gha16, LRY17, GU19, Gha22], maximal matching [MV13, YYI09, LRY17, BRR23], and $(\Delta+1)$ vertex coloring [EMR14, FPSV17, CMV18, ${}^{+}$19]. For other problems, polylogarithmic complexities are ruled out by lower bounds. For example, given a graph $G$, for the task of providing local access to a spanner of $G$, the best known query complexities [LRR16, LL18, PRVY19, LRR20, ACLP23] are $\widetilde{O}(n^{2/3})$, and known lower bounds imply that $\Omega(\sqrt{n})$ time is required even for constant degree graphs [LRR16].
1.1 Our Results: Average-Case LCAs
@@
We say that $\mathcal{A}$ is an average-case local computation algorithm for a distribution over objects $\mathcal{G}$ if, with probability $(1-1/n)$ over $G\leftarrow\mathcal{G}$, $\mathcal{A}^{G}$ is an LCA.
Note that our definition requires that with high probability over the object $G$, the LCA succeeds for every query on this object.
We consider average-case LCAs for two well-studied distributions over graphs: the Erdős-Rényi and Preferential Attachment models. To demonstrate the large advantage that average-case assumptions can provide, we focus on the well-studied problem of LCAs for spanners [LRR16, LL18, LRR20, PRVY19, ACLP23, BF24].
@@
For general graphs, a spanner with size $O(n^{1+1/k})$ and stretch $(2k+1)$ can be constructed in linear time by a centralized algorithm [BS03]. Moreover, conditional on Erdos’ girth conjecture [Erd64] this size-stretch tradeoff is tight.
@@
In fact, for general graphs there is a lower bound of $\Omega(\sqrt{n})$ work per query [LRR16, PRVY19], even for graphs of bounded degree. In fact, this lower bound applies for the weaker problem of providing access to a sparse connected subgraph (with no stretch requirement). Thus, to obtain faster algorithms we must make a “beyond worst case” assumption. For both random graph models we consider, our algorithms achieve a size-stretch tradeoff that is impossible to achieve for general graphs under Erdos’ girth conjecture, while simultaneously achieving a query time that is impossible for LCAs for general graphs.
1.1.1 Local Computation Algorithms for the Erdős-Rényi Model.
@@
To motivate our results, we first overview a very simple construction that we will compare against. For a graph $G\leftarrow G(n,p)$, if we keep each edge in $G$ with probability $p^{\prime}/p$ for some $p^{\prime}<p$, we effectively sample a graph $H\subseteq G$ that is itself distributed as $G(n,p^{\prime})$. It is well known that for any $p^{\prime}\geq p_{0}=(2+\varepsilon)\log(n)/n$, this graph will be connected with high probability. Moreover, the graph will be an expander whp and hence will have diameter (and thus stretch, when considered as a spanner of $G$) $O(\log n)$. It is immediate to implement an LCA that keeps each edge of $G$ with probability $\min\{p_{0}/p,1\}$ [^1] The LCA uses its random tape to answer future queries to the same edge in a consistent fashion., and we can summarize the resulting algorithm in the following:
Observation 1.4.
@@
This algorithm improves the edge count of Observation 1.4, but retains superconstant stretch (and has a slower query time). We obtain results that improve on both constructions, obtaining ultra-sparse spanners (i.e. with $n+o(n)$ edges) and constant stretch, while maintaining fast query time.
Our first LCA considers random graphs with relatively large degree, and improves both the stretch and the sparsity of Observations 1.4 and 1.5. If we assume the adjacency list of each vertex is sorted, the LCA has constant probe complexity. If this is not the case, the LCA has probe complexity bounded by $\min\{\Delta,\log(n)/p\}$ whp. In particular, for highly dense and highly sparse graphs, we obtain a runtime $n^{\varepsilon}$ for small $\varepsilon$, beating the worst-case lower bound of $\Omega(\sqrt{n})$.
@@
For every $np=n^{\delta}$, there is an average-case LCA for $G\leftarrow G(n,p)$ that whp gives access to a $(2/\delta+5)$-spanner with $n+o(n)$ edges. Moreover, the LCA has probe complexity $1$ where we have access to a sorted adjacency list in $G$, and $O\left(\min\left\{n^{\delta},n^{1-\delta}\log n\right\}\right)$ otherwise.
A slight modification of our result allows constant query complexity even without sorted adjacency lists, at the cost of including a nearly-linear number of edges.
Theorem 1.7.
For every $np=n^{\delta}$, there is an average-case LCA for $G\leftarrow G(n,p)$ that whp gives access to a $(2/\delta+5)$-spanner with $O(n\log n)$ edges. Moreover, the LCA has probe complexity $1$.
This result is superior to Observation 1.4 and incomparable to Observation 1.5, as the stretch remains constant.
Finally, we show that for sufficiently dense graphs, we can provide an alternate construction with even better (constant) stretch, with a nearly matching query time:
Theorem 1.8.
For every $np=n^{\delta}$ for $\delta>1/2$, there is an average-case LCA for $G\leftarrow G(n,p)$ that whp gives access to a $4$-spanner with $n+o(n)$ edges. Moreover, the LCA has probe complexity $\widetilde{O}(n^{\delta})$.
Again we improve both the stretch and size of Observations 1.4 and 1.5.
Our last result in the Erdős-Rényi model focuses on sparse input graphs (for instance, those with $np=n^{o(1)}$.) Here we consider the relaxed problem of producing a sparse spanning subgraph (LCAs for which have been studied before [${}^{+}$15, LL18, PRVY19, LRR20, BF24]), where we do not bound the stretch. We note that known lower bounds [${}^{+}$15, PRVY19] imply a $\sqrt{n}$ query lower bound even for this problem.
We are able to obtain an ultra-sparse connectivity-preserving subgraph for all edge probabilities greater than $p^{*}=7\log(n)/n$, only a constant factor above the connectivity threshold. Moreover, we achieve query time $\widetilde{O}(\Delta)$, where $\Delta=np$ is the expected degree of the graph (compared to the worst-case lower bound of $\Omega(\sqrt{n})$).
@@
1.1.2 Local Computation Algorithms for the Preferential Attachment Model.
In the preferential attachment model (formally defined in Definition 2.2), we construct a graph by sequentially inserting $n$ vertices. For each new vertex $v_{i}$, we add $\mu=\mu(n)$ edges from $v_{i}$ to existing vertices $(v_{1},\ldots,v_{i-1})$, where we add an edge to $v_{j}$ with probability proportional to the degree of $j$. (Afterwards, the vertices are permuted randomly, so that the algorithm cannot use the IDs to determine insertion order). Such a model captures the property that high-degree nodes tend to accumulate additional connections. The generation of preferential attachment graphs (and variants of it) have been extensively studied [AKM13, BB05, ${}^{+}$00, MP16, NLKB11, YH10].
We construct spanner LCAs for preferential attachment graphs of sufficiently high degree. In fact, our LCA constructs a stronger object - a low stretch spanning tree. A low-stretch spanning tree $H\subseteq G$ is a spanner with exactly $n-1$ edges, the minimum required even to preserve connectivity. Such an object (with nontrivial stretch) clearly does not exist for general graphs, but we construct (via an LCA) a spanning tree with $O(\log n)$ stretch.
@@
For every preferential attachment process with parameter $\mu>c_{\mu}\log n$ for a global constant $c_{\mu}$, there is an average-case LCA for $G\leftarrow G_{pa}(n,\mu)$ that w.h.p gives access to an $O(\log n)$-spanner $H\subseteq G$. Moreover:
  • on query $(u,v)$ the LCA has time complexity $O(d_{u}+d_{v})$, which is $O(\sqrt{n})$ in the worst-case and $O(\log^{3}n)$ in expectation (over all possible queries),
  • $H$ contains $n-1$ edges.
A Note on The Access Model.
In several recent works on LCAs for spanners, there has been a question of whether vertices are given sequential IDs from $\{1,\ldots,n\}$, or an ID from a larger global universe (say, $[n^{10}]$). Equivalently, is the LCA allowed to probe vertices that it has not discovered by local search? The LCAs of Theorem 1.10 and Theorems 1.7, 1.9 and 1.8 work under the weaker assumption that vertex IDs are drawn from an arbitrary polynomial sized universe. Theorem 1.6 does require this stronger model, but we stress that this extra power is a standard feature of the LCA model. To support this claim, we refer the reader to the formal definition in [Gol17] (Definition 12.11) as well as [AD, EMR18]. This probe model is utilized in frontier work on graph coloring and approximate maximum matching [Gol17, LRY17, BRR23].
1.2 Our Results: Joint Sampling
Finally, a natural topic relating to local algorithms and random graphs is to sample the random graph itself in a local fashion, rather than assuming we have probe access to one that already exists. Several recent works [GGN03, AN08, BRY20, BPR22, MSW22, ELMR21] studied exactly this question, under the label of Local Access Generators (LAGs). These algorithms provide efficient query access to a random instance of some structure.

### v2

# Beyond Worst Case Local Computation Algorithms

@@
Beyond Worst Case Local Computation Algorithms
Amartya Shankha Biswas
MIT
asbiswas@mit.edu
Supported by NSF award CCF-2310818.
   Ruidi Cao
MIT
ruidicao@mit.edu
   Cassandra Marcussen
Harvard University
cmarcussen@g.harvard.edu
Supported in part by an NDSEG fellowship, and by NSF Award 2152413 and a Simons Investigator Award to Madhu Sudan.
   Edward Pyne
MIT
epyne@mit.edu
Supported by a Jane Street Graduate Research Fellowship and NSF awards CCF-2310818 and CCF-2127597.
   Ronitt Rubinfeld
MIT
ronitt@csail.mit.edu
Supported by NSF awards CCF-2006664, DMS-2022448 and CCF-2310818.
   Asaf Shapira
Tel Aviv University
asafico@tauex.tau.ac.il
Supported in part by ERC Consolidator Grant 863438.
   Shlomo Tauber
Tel Aviv University
shlomotauber@mail.tau.ac.il
Supported in part by ERC Consolidator Grant 863438.
We initiate the study of Local Computation Algorithms on average case inputs. In the Local Computation Algorithm (LCA) model, we are given probe access to a huge graph, and asked to answer membership queries about some combinatorial structure on the graph, answering each query with sublinear work. We define a natural model of average-case local computation algorithms.
For instance, an LCA for the $k$-spanner problem gives access to a sparse subgraph $H\subseteq G$ that preserves distances up to a multiplicative factor of $k$. Our first result builds LCAs for this problem assuming the input graph is drawn from a variety of well-studied random graph models – Preferential Attachment, Uniform Attachment, and Erdős-Rényi with a variety of parameters. Our spanners achieve size and stretch tradeoffs that are impossible to achieve for general graphs, while having dramatically lower query complexity than known worst-case LCAs.
Finally, we investigate the intersection of LCAs with Local Access Generators (LAGs). Local Access Generators provide efficient query access to a random object. We explore the natural problem of generating an Erdős-Rényi random graph together with a combinatorial structure on it. We show that this combination can be easier to solve than focusing on each problem by itself, by building a fast, simple algorithm that provides access to an Erdős-Rényi random graph together with a maximal independent set.
1 Introduction and Our Results
When computing on a very large object, it can be important to find fast algorithms which answer user queries to the solution, while neither considering the whole input, nor computing the full output solution. In the local computation model [RTVX11, ARVX12], we are given probe access to a large object, such as a graph, and receive queries about some combinatorial structure on the graph. We desire Local Computation Algorithms (LCAs) that can quickly answer such queries while making very few probes to the graph. Moreover, we require the answers returned by the algorithm on different queries to be consistent with a fixed global structure. This consistency requirement is challenging since typically, we require the LCA to be memoryless – it does not store information about its previous answers.
@@ 1 Introduction and Our Results
A Local Computation Algorithm (LCA) for a problem $\Pi$ is an oracle $\mathcal{A}$ with the following properties. $\mathcal{A}$ is given probe access to input $G$, a sequence of random bits $\vec{r}$ and a local memory. For any query $q$ in a family of admissible queries to the output, $\mathcal{A}$ must use only its oracle access to $G$ (which we refer to as probes to $G$), random bits $\vec{r}$, and local memory to answer the query $q$. After answering the query, $\mathcal{A}$ erases its local memory (including the query $q$ and its response). Let $T_{\mathcal{A}}(G,q)$ denote the expected (over the choice of $\vec{r}$) number of probes it takes for the LCA $\mathcal{A}$ to answer query $q$ on input $G$, and set $T_{\mathcal{A}}(G)=\max_{q}T_{\mathcal{A}}(G,q)$. We say the LCA has probe complexity $T(n)$ if the maximum of $T_{\mathcal{A}}(G)$ over all possible inputs $G$ parametrized by size $n$ is $T(n)$. All the responses to queries given by $\mathcal{A}$ must be consistent with a single valid solution $X$ to the specified computation problem on input $G$.
There has been extensive work on fast LCAs for a variety of natural problems. For example, on bounded degree graphs, there are LCAs with polylogarithmic query complexity for maximal independent set (MIS) [Gha16, LRY17, GU19, Gha22], maximal matching [MV13, YYI09, LRY17, BRR23], and $(\Delta+1)$ vertex coloring [EMR14, FPSV17, CMV18, CFG+19]. LCAs have found applications in well-studied algorithmic problems [ABGR25, LLRV25a] (such as matching) and have contributed to breakthroughs in learning theory [LLRV25b, LV25a, LV25b].
For other problems, polylogarithmic complexities for LCAs are ruled out by lower bounds. For example, given a graph $G$, for the task of providing local access to a spanner of $G$, the best known query complexities [LRR16, LL18, PRVY19, LRR20, ACLP23, BF24] are $\widetilde{O}(n^{2/3})$, and known lower bounds imply that $\Omega(\sqrt{n})$ time is required even for constant degree graphs [LRR16].
@@
We say that $\mathcal{A}$ is an average-case local computation algorithm for a distribution over objects $\mathcal{G}$ parametrized by size $n$ for problem $\Pi$ if, with probability $(1-1/n)$ over $G\leftarrow\mathcal{G}$, $\mathcal{A}^{G}$ (the algorithm when given probe access to input $G$) is an LCA. We say that the LCA $\mathcal{A}$ has average-case probe complexity $T(n)$ if the expected probe complexity $T_{\mathcal{A}}(G)$ over $G\leftarrow\mathcal{G}$ is $T(n)$. We say the LCA has worst-case probe complexity $T(n)$ if the maximum probe complexity $T_{\mathcal{A}}(G)$ over $G\leftarrow\mathcal{G}$ is $T(n)$.
Note the requirement that with high probability over the object $G$, the LCA succeeds for every query on this object.
Remark 1.3.
[…]
We now describe our results on average case LCAs for graph spanners. We then describe a new model of local access generation which locally generates a random object together with a combinatorial structure, and give our results for locally generating a random graph together with a maximal independent set.
1.1 Our Results: Spanner LCAs for Average-Case Graphs
Our first set of results focus on the well-studied problem of LCAs for spanners [LRR16, LL18, LRR20, PRVY19, ACLP23, BF24]. We study LCAs for spanners over the Erdős-Rényi, Preferential Attachment, and Uniform Attachment random graph models.
@@ 1.1 Our Results: Spanner LCAs for Average-Case Graphs
For general graphs, a spanner with size $O(n^{1+1/k})$ and stretch $(2k+1)$ can be constructed in linear time [BS03]. Moreover, conditional on Erdos’ girth conjecture [Erd64] this size-stretch tradeoff is tight.
@@
For general graphs there is a lower bound of $\Omega(\sqrt{n})$ work per query [LRR16, PRVY19], even for graphs of bounded degree.
Thus, to obtain faster algorithms we must make a “beyond worst case” assumption. For the random graph models we consider, our algorithms achieve a size-stretch tradeoff that is impossible to achieve for general graphs under Erdos’ girth conjecture, while simultaneously achieving a query time that is impossible for LCAs for general graphs.
1.1.1 Local Computation Algorithms for the Erdős-Rényi Model
@@
To motivate our results, we first overview two simple constructions that we will compare against. First, for a graph $G\leftarrow G(n,p)$, if we keep each edge in $G$ with probability $p^{\prime}/p$ for some $p^{\prime}<p$, we effectively sample a graph $H\subseteq G$ that is itself distributed as $G(n,p^{\prime})$. It is well known that for any $p^{\prime}\geq p_{0}=(2+\varepsilon)\log(n)/n$, this graph will be connected with high probability. Moreover, the graph will be an expander whp and hence will have diameter (and thus stretch, when considered as a spanner of $G$) of $O(\log n)$. It is immediate that we can implement an LCA that keeps each edge of $G$ with probability $\min\{p_{0}/p,1\}$ [^1] The LCA uses its random tape to answer future queries to the same edge in a consistent fashion., and we can summarize the resulting algorithm in the following:
Observation 1.5.
@@
This algorithm improves the edge count of Observation 1.5, but retains superconstant stretch (and has a slower query time). We improve on both constructions, by obtaining ultra-sparse spanners (i.e. with $n+o(n)$ edges) and constant stretch. For dense graphs, our results are as follows:
@@
For every $np=n^{\delta}$, there is an average-case LCA for $G\leftarrow G(n,p)$ that whp gives access to a $(2/\delta+5)$-spanner $H$ with $n+o(n)$ edges. Moreover, the LCA has probe complexity $1$ where we have access to a sorted adjacency list in $G$, and $O\left(\min\left\{n^{\delta},n^{1-\delta}\log n\right\}\right)$ otherwise.
In particular, for highly dense and highly sparse graphs, we obtain a runtime $n^{\varepsilon}$ for small $\varepsilon$, beating the worst-case lower bound of $\Omega(\sqrt{n})$.
Our last result in the Erdős-Rényi model focuses on sparse input graphs (for instance, those with $np=n^{o(1)}$.) Here we consider the relaxed problem of producing a sparse spanning subgraph (LCAs for which have been studied before [LMR+17, LL18, PRVY19, LRR20, BF24]), where we do not bound the stretch. We note that known lower bounds [LMR+17, PRVY19] imply a $\sqrt{n}$ query lower bound even for this problem on sparse graphs.
We are able to obtain an ultra-sparse connectivity-preserving subgraph for all edge probabilities greater than $p^{*}=7\log(n)/n$, only a constant factor above the connectivity threshold. Moreover, we achieve query time $\widetilde{O}(\Delta)$, where $\Delta=np$ is the expected degree of the graph.
@@
1.1.2 Local Computation Algorithms for the Preferential and Uniform Attachment Models
Next, we construct spanner LCAs for preferential and uniform attachment graphs with a sufficiently high degree parameter. In the preferential attachment model (formally defined in Definition 6.1), the graph is constructed by sequentially inserting $n$ vertices. For each new vertex $v_{i}$, $\mu=\mu(n)$ edges are added to the graph; these may either be self loops or edges from $v_{i}$ to existing vertices $(v_{1},\ldots,v_{i-1})$, where an edge is added to $v_{j}$ with probability proportional to the degree of $j$. (Afterwards, the vertices are permuted randomly, so that the algorithm cannot use the IDs to determine insertion order). Such a model captures the property that high-degree nodes tend to accumulate additional connections. The generation of preferential attachment graphs (and variants of it) have been extensively studied [AKM13, BB05, KRR+00, MP16, NLKB11, YH10].
Our spanner LCA for preferential attachment graphs of sufficiently high degree constructs a low stretch spanning tree, a stronger object than a spanning subgraph. A low-stretch spanning tree $H\subseteq G$ is a spanner with exactly $n-1$ edges, the minimum required even to preserve connectivity.
@@ 1.1.2 Local Computation Algorithms for the Preferential and Uniform Attachment Models
For every preferential attachment process with parameter $\mu>c_{\mu}\log(n)$ for a global constant $c_{\mu}$, there is an average-case LCA for $G\leftarrow G_{pa}(n,\mu)$ that w.h.p gives access to an $O(\log n)$-spanner $H\subseteq G$, and moreover $H$ contains $n-1$ edges. On query $(u,v)$ the LCA has time complexity $O(d_{u}+d_{v})$, which is $O(\mu\sqrt{n})$ in the worst-case and $O(\mu\log^{3}n)$ in expectation (over all possible queries).
We note that techniques from previous work on local information algorithms [FP14] can be applied to obtain an LCA giving access to a $\tilde{O}(\log n)$-spanner with $\tilde{O}(n)$ edges with a query time that is $O(\mu\sqrt{n})$ in the worst case and $O(\mu\operatorname{polylog}(n))$ in expectation for any $\mu$ (see Section 6). We focus on the setting where $\mu$ is sufficiently large and the resulting graph is therefore not already sparse. Our LCA in this setting achieves improved bounds for the sparsity and query time.
We also construct spanner LCAs for uniform attachment graphs of sufficiently high degree. In the uniform attachment model (formally defined in Definition 7.1), $n$ vertices are also inserted sequentially. At each time step, a new vertex $v_{i}$ joins the graph and $\mu=\mu(n)$ edges are added from $v_{i}$ to existing vertices $(v_{1},\ldots,v_{i-1})$, which are each chosen independently and uniformly at random. We assume again that the insertion order of vertices is unknown to the LCA. Uniform attachment graphs [TM67] are standard models of random circuits and randomly evolving networks with applications including the modeling of networks, physical processes, and the spread of contamination among organisms, as noted in [ZZ15].
Theorem 1.10.
For every uniform attachment process with parameter $\mu>c_{\mu}\log^{2}(n)$ for a global constant $c_{\mu}$, there is an average-case LCA for $G\leftarrow G_{ua}(n,\mu)$ that w.h.p gives access to an $O(\log n)$-spanner $H\subseteq G$, and moreover $H$ contains $n+c$ edges, for some constant $c$ independent of $n$.
Moreover, let $D:=\mu\cdot(H_{n-1}-H_{6})+\mu/2$ where $H_{n}$ denotes the $n$-th Harmonic number. On query $(u,v)$, if $\min\{d_{u},d_{v}\}>D$, the time complexity is $O(1)$. Otherwise, the time complexity is $O(d_{u}+d_{v})$ which is $O(\mu\log n)$ in the worst-case and $O(\mu)$ in expectation (over all possible queries).
1.2 Our Results: Joint Sampling of Erdős-Rényi Graphs and Maximal Independent Sets
A natural topic relating to local algorithms and random graphs is to sample the random graph itself in a local fashion, rather than assuming we have probe access to one that already exists. Several recent works [GGN03, AN08, BRY20, BPR22, MSW22, ELMR21] studied exactly this question, under the label of Local Access Generators (LAGs). These algorithms provide efficient query access to a random instance of some structure.